A 0.145- kg baseball with an initial speed of 29 m/s rebounds with a speed of 39 m/s after being struck with a bat. If the duration of contact between ball and bat was 1.6 ms, what was the magnitude of the average force between ball and bat?
Answer #1

u=29m/s , v=39m/s ,m=0.145Kg , t=1.6s
v=u+at
39=29+a*1.6
a=10/1.6
F=ma
=0.145*10/1.6.1=0.90N
Answer #3

magnitude of average force is

F(avg) = dp/dt = m(v2-v1)/t = 0.145 (39-29)/1.6*10^-3 s =906.25 N